A bike is accelerating in yz-plane with its speed given by ()


at


t(w2) +2 * u2: dvt)


at


= u2


= u3 * t(2+w3)+3*u3, subject to the initial conditions,


v,(0) = n2; v0) = n3. Determine its speed at later time


u2 = 4 + 0.3R, u3 = 2+ 0.4R, w2 = 0.5+ R,w3 = 1+ R, n2 = 3.1+ 0.2R, n3 = 4.1 + 0.1R, p1 = 5.4 + 0.2R


1
Expert's answer
2021-04-15T07:39:12-0400

"\frac{\partial v_y(t)}{\partial t}=u_2\cdot3t^{w_2},\ \frac{\partial v_z(t)}{\partial t}=u_3\cdot t^{w_3}+3"

"v_y(0)=n_2,\ v_z(0)=n_3"


"v_y(t)=\frac{3u_2t^{w_2+1}}{w_2+1}+n_2" , "v_z(t)=\frac{u_3t^{w_3+1}}{w_3+1}+n_3"


"v_y(p_1)=\frac{3(3+0.3R)(5.4+0.2R)^{2+0.3R}}{2+0.3R}+3.1+0.2R"


"v_z(p_1)=\frac{(5+0.4R)(5.4+0.2R)^{2+R}}{2+R}+4.1+0.1R"